# MotionAgent: Fine-grained Controllable Video Generation via Motion Field Agent

### v1

@@ Abstract
We propose MotionAgent, enabling fine-grained motion control for text-guided image-to-video generation. The key technique is the motion field agent that converts motion information in text prompts into explicit motion fields, providing flexible and precise motion guidance. Specifically, the agent extracts the object movement and camera motion described in the text and converts them into object trajectories and camera extrinsics, respectively. An analytical optical flow composition module integrates these motion representations in 3D space and projects them into a unified optical flow. An optical flow adapter takes the flow to control the base image-to-video diffusion model for generating fine-grained controlled videos. The significant improvement in the Video-Text Camera Motion metrics on VBench indicates that our method achieves precise control over camera motion. We construct a subset of VBench to evaluate the alignment of motion information in the text and the generated video, outperforming other advanced models on motion generation accuracy.
@@ 1 Introduction
Some I2V generation models [12, 4] randomly make the input image dynamic, resulting in an uncontrollable generated result. Other I2V models [54, 68, 37, 9, 66, 29, 30] support simple control through text input. As shown in Figure 1(a), they adopt a text encoder to inject control information into visual features. They can perform overall control, but it is difficult to achieve fine-grained control over each element of the video. Additionally, training an I2V generation model with high alignment between motion information in the text and the video relies heavily on the quality of training data [3, 58, 49, 7, 42], and obtaining high-quality training data is also labor-intensive.
@@
As illustrated in Figure 1(b), some controllable I2V generation models adopt specially designed modules for controlling different motion types, such as object movement [50, 61, 52, 71, 69, 32, 51] or camera motion [14, 55, 56]. These modules are proposed for individual control conditions, making it difficult to control multiple motion types simultaneously. Additionally, inputting these control conditions may require expert knowledge, which creates a barrier for non-professional users. These methods are less flexible than those that directly control video motion through text input.
@@
To this end, we propose MotionAgent, which enables fine-grained control for text-guided I2V generation. The key design is the motion field agent that converts the motion information in the text prompts into explicit motion representations, providing precise motion guidance. As shown in Figure 1(c), our framework directly takes the text as input. First, the agent parses and converts the motion information in the text into object trajectories and camera extrinsics, which explicitly represent object movement and camera motion, respectively. Then, these two intermediate representations are integrated in the 3D space and projected into unified optical flow maps through the proposed analytical optical flow composition module. Finally, we tune an optical flow adapter, using the unified flow as conditions, to control a base I2V diffusion model for video generation.
@@
In the experiments, we first evaluate our method on a public I2V generation benchmark [20]. Evaluation results demonstrate that our approach significantly improves the control accuracy of camera motion and achieves comparable video quality with other advanced models. Furthermore, we construct a subset of VBench for assessing the alignment of motion information in the text and the generated video. The results illustrate that our method achieves the best motion accuracy. We also conduct a user study, which shows that our generated videos align better with the motion information in the text and maintain high quality. Finally, we conduct ablation studies to verify the effectiveness of the proposed components. Our contributions can be summarized as follows:
@@ 2.1 Video Generation via Agent
The development of ChatGPT [34] and GPT-4 [1] represents a crucial advancement in natural language processing. These large language models (LLMs) [48, 63, 43, 26, 67] enable multi-round conversations and have the impressive ability to follow complex instructions. Furthermore, the integration of vision capabilities in GPT-4V [59] is a milestone, enabling LLMs to process and interpret multimodal data. The vision capabilities in LLMs pave the way for the development of multimodal LLM-powered agents. These agents increasingly gain attention for their ability to effectively handle multimodal inputs and achieve complex tasks.
A series of studies propose to exploit the knowledge of agents for achieving controllable generation [24, 10, 25], zero-shot generation [19, 28, 18, 33], or long video generation [72, 47]. Anim-Director [22] builds an autonomous animation-making agent for generating contextually coherent animation videos. DreamFactory [53] leverages multi-agent collaboration principles and a keyframe iteration method to ensure consistency and style across long videos. VIDEOAGENT [41] aims to improve video plan generation based on a feedback pipeline. ChatCam [27] introduces a system that navigates camera motion through conversations with users and mimics a professional cinematographer’s workflow.
Our research contributes to this area by proposing a motion field agent to decompose and convert the motion information in the text into explicit intermediate representations of object movement and camera motion to guide precise control over video generation.
@@ 2.2 Text-guided I2V Generation
Generally, most I2V generation models [9, 66, 29, 21] support text and image input simultaneously. The input image guides the visual content of the video, while the text indicates the potential motion. VideoCrafter [6], Dynamicrafter [54], ConsistI2V [37], and I2VGen-XL [68] are all based on the U-Net [38] backbone and use a text encoder to inject text embeddings into visual features, enabling text control over video content. CogVideoX [60] and Open-Sora [45], based on the DIT [35] architecture, gain control capability by jointly learning text tokens and visual tokens. These methods, which directly learn the similarity between text embeddings and visual features, often highly depend on the quality of training data. However, there is a lack of large high-quality datasets with text labels that describe video motion. The text labels in most video datasets [3, 58, 49, 7, 42] describe the content of a frame without much motion information and are not suitable for training I2V generation models with strong motion control capability, which leads to poor performance of these methods in achieving precise motion control through direct text input.
@@ 2.3 I2V Generation via Motion Control Module
With the development of video generation models, controllable video generation gradually attracts more attention. Existing methods [69, 55, 56] propose designing specialized modules for certain motion types, such as object movement or camera motion. DragNUWA [61] encodes sparse point trajectories into dense features as guidance information, which is then injected into the diffusion model to control object motion. DragAnything [52] uses masks to identify a central point and subsequently generates a Gaussian map to track this center, providing a guiding trajectory for object-controllable I2V generation. TrackGo [71] introduces a TrackAdapter, which encodes object trajectories into the network through a dual attention mechanism to control object movement in the generated videos. Other video generation methods achieve control of camera motion. MotionCtrl [50] directly encodes camera extrinsics as features and inputs them to the model. IMAGE CONDUCTOR [23] and CameraCtrl [14] both convert sparse camera extrinsics into dense feature maps using Plücker embeddings and employ these dense embeddings to control camera motion.
In addition to encoding control conditions into deep features, some methods generate explicit intermediate representations, such as optical flow, to control the generated video. Motion-I2V [39] factorizes I2V generation into two stages: It first generates optical flow maps using a flow diffusion model and then uses the optical flow maps as explicit motion guidance for video generation. MOFA-Video [32] proposes a generative motion field adapter. They first convert sparse point trajectories into dense optical flow through a sparse-to-dense network and then use an optical flow adapter to control video generation. Although these models achieve control over the videos via a motion control module, and their control capabilities are stronger than those of text-based models, they often require specific knowledge, such as providing object trajectories or camera extrinsics, which creates a barrier for users. Moreover, most controllable I2V generation methods can only control a single motion type at a time and struggle to achieve simultaneous control of all types.
Our method leverages the optical flow adapter proposed by MOFA-Video [32] to control the base I2V diffusion model. Unlike MOFA-Video, our approach has two differences: 1) We propose a motion field agent that allows direct control of the generated video through text input rather than manual trajectory plotting. 2) We introduce an analytical optical flow composition module, which enables our method to simultaneously control both object movement and camera motion.
@@ 3 Method
Figure 2: Pipeline of Motion Field Agent. Step 1: The agent first parses the input text, dividing the motion information into two parts that respectively describe object movement and camera motion. Step 2: The agent draws the object trajectories according to the text of object movement. Step 3: The agent directly generates the camera extrinsics based on the text of camera motion.
Our method aims to achieve fine-grained controllable I2V generation through text input, allowing precise control of both object movement and camera motion. Specifically, a motion field agent first converts the motion information in the text into object trajectories and camera extrinsics. These two explicit intermediate representations are then fed into an improved controllable I2V generation model to guide the video generation. Within the controllable I2V generation model, the object trajectories and camera extrinsics are integrated by an analytical optical flow composition module to compute unified optical flow maps. Subsequently, we employ Stable Video Diffusion (SVD) [4] as the base I2V diffusion model and utilize an optical flow adapter [32] as the motion control module to generate the final video.
@@ 3.1 Motion Field Agent
As for the motion field of a video, we generally decompose it into object movement within the frame and overall camera motion. Based on the above decomposition, our agent performs two key tasks: object trajectory plotting and camera extrinsics generation. The details of the designed agent can be found in Figure 2.
@@ 3.1.2 Object Trajectory Plotting
Trajectory Plotting: (4) The agent locates the object in the image based on the detection and segmentation results, thereby determining the starting point of the trajectory. (5) The agent plots the trajectory with varying lengths and curvature according to the complexity of the text description.
To enhance the capabilities of trajectory plotting, we replace direct trajectory generation with a grid selection approach [65]. An image is divided into $N\times M$ grids, with each grid labeled by an integer. The agent determines each point of the trajectory by selecting grid numbers, and the trajectory is formed by sequentially connecting all the points. More details of trajectory plotting can be found in the supplementary material.
@@ 3.1.3 Camera Extrinsics Generation
Agent directly generates camera extrinsics $E$ based on the text of camera motion and the input image. The text description specifies the motion type, such as zoom in/out, pan left/right, and tilt up/down. The image helps the agent determine the magnitude of the camera motion. For instance, a wide landscape scene may require intensive camera motion, while a narrow close-up shot may only need slight adjustments of camera location. It is worth noting that we constrain the translation $T$ of the camera extrinsics to $(-1,1)$. In the next analytical optical flow composition module, we rescale the translation $T$ back to a reasonable range based on the estimated depth map.
@@ 3.2.1 Analytical Optical Flow Composition
We first use Metric3D [62] to estimate the depth map $D$ of the input image and unproject each pixel $I^{0}$ in the image to obtain the 3D location $P$.
CMP [64] is utilized to estimate the dense optical flow of object movement $F_{obj}$ according to the sparse object trajectories. Based on the estimated optical flow $F_{obj}$ and the depth map $D$, we calculate the 3D location offset $O$ caused by object movement. Note that we assume a large focal length for all images, thereby neglecting the offset on the z-axis, where $o_{z}=0$. We simply assume that the 3D location offset of the $i$-th pixel on the x/y-axis $o^{i}_{xy}$ is proportional to its depth $d^{i}$, with the same direction as the corresponding optical flow $f^{i}_{obj}$. It can be formulated as,
$o^{i}_{xy}=K^{-1}\cdot{f^{i}_{obj}}\cdot{d^{i}},$ (1)
where $K$ is the camera intrinsics.
Then, we apply the offset on the x/y-axis $o^{i}_{xy}$ to the 3D location of $i$-th pixel $p^{I}_{xy}$, which can be computed as,
$\hat{p}^{i}_{xy}=p^{I}_{xy}+o^{i}_{xy},$ (2)
where $\hat{p}^{i}_{xy}$ is the new 3D location of $i$-th pixel on the x/y-axis, and $\hat{p}^{i}_{z}=p^{i}_{z}$.
We rescale the translation $T$ of the camera extrinsics $E$ according to the maximum depth range. Subsequently, we reproject these moved points $\hat{P}$ into the corresponding image coordinate systems based on the camera extrinsics $E^{t}$ to calculate the new pixel position $I^{t}$ in the $t$-th frame. It can be calculated as,
$I^{t}=\Pi(E^{t}\hat{P}),$ (3)
where, $E^{t}$ is the camera extrinsics of the $t$-th frame, and $\Pi$ is the projection operation. We assume that the camera coordinate of the first frame serves as the world coordinate.
Finally, we calculate the pixel offsets of the corresponding pixel points, which can be noted as,
$\bar{F}^{t}=I^{t}-I^{0},$ (4)
where $\bar{F}$ are named unified optical flow maps. These flow maps contain motion information of both object movement and camera motion.
@@ 3.2.2 Optical Flow Adapter Tuning
Since the unified optical flow maps involve many simplifications and approximations, they contain geometrically unrealistic regions compared to the true optical flow maps. Directly using these flow maps as input for the optical flow adapter, which is trained on real optical flow maps, may lead to degraded results in the generated videos. We propose fine-tuning the optical flow adapter on the unified optical flow maps.
For each training video, we first use an optical flow model, Unimatch [57], to estimate the real optical flow. Then we employ a SLAM method, DROID-SLAM [46], to compute the camera extrinsics for each frame. Next, we eliminate the optical flow caused by camera motion based on the estimated camera extrinsics, and approximately obtain the optical flow caused by object movement. We perform sparse sampling [64] on these optical flow maps to obtain sparse object trajectories. More details of data preparation can be found in the supplementary material. Subsequently, we reuse the proposed analytical composition method to calculate the unified optical flow maps, which are utilized as input to fine-tune the optical flow adapter. This step eliminates the domain gap between real and unified flow fields and facilitates high-quality video results.
@@ 4 Experiments
We provide extensive comparisons to evaluate our models on controllable I2V generation tasks. First, we introduce the implementation details and evaluation metrics (Sec. 4.1). Importantly, we show the results on the general I2V generation benchmark (Sec. 4.2.1), controllable I2V generation benchmark (Sec. 4.2.2), and user study (Sec. 4.2.3). Lastly, we conduct ablation studies to demonstrate the effectiveness of each module in our MotionAgent (Sec. 6.1).
@@ 4.1 Implementation Details
Metrics. In the comparison experiments of the general I2V generation task, we adopt the same evaluation metrics introduced by VBench [20]. In the comparison of controllable I2V generation, we report Object Movement Q&A, Complex Camera Motion, and Overall Score.
I2V Score reports the overall score of I2V generation metrics. Video-Text Camera Motion assesses the consistency between camera motion and the input text, such as zoom in/out. Video-Image Subject Consistency assesses whether the appearance of the subject remains consistent throughout the entire video compared to the input image. Video-Image Background Consistency evaluates the temporal consistency of background scenes with the input image. Subject Consistency assesses whether the subject’s appearance remains consistent throughout the entire video. Background Consistency evaluates the temporal consistency of the background scenes across frames. Motion Smoothness evaluates whether the motion in the generated video is smooth and follows the physical laws of the real world. Aesthetic Quality evaluates the artistic and aesthetic value perceived by humans towards each video frame. Dynamic Degree evaluates the level of dynamics generated by each model. Object Movement Q&A assesses the consistency between the text description and object movement in the video. Complex Camera Motion evaluates the consistency between complex camera movements in the generated video and the input text description. Total Score is the overall metric of controllable I2V generation. All evaluation metrics are higher-the-better.

### v2

@@ Abstract
We propose MotionAgent, enabling fine-grained motion control for text-guided image-to-video generation. The key technique is the motion field agent that converts motion information in text prompts into explicit motion fields, providing flexible and precise motion guidance. Specifically, the agent extracts the object movement and camera motion described in the text, and converts them into object trajectories and camera extrinsics, respectively. An analytical optical flow composition module integrates these motion representations in 3D space and projects them into a unified optical flow. An optical flow adapter takes the flow to control the base image-to-video diffusion model for generating fine-grained controlled videos. After that, an optional rethinking step can be adopted to ensure the generated video is aligned well with motion information in the prompt. The significant improvement in the Video-Text Camera Motion metrics on VBench indicates that our method achieves precise control over camera motion. We further construct a subset of VBench to evaluate the alignment of motion information in the text and the generated video, outperforming other advanced models on motion generation accuracy.
@@ 1 Introduction
Some I2V generation models [18, 5] randomly make the input image dynamic, resulting in an uncontrollable generated result. Other I2V models [63, 74, 44, 11, 73, 36, 37] support simple control through text input. As shown in Figure 1(a), they adopt a text encoder to inject control information into visual features. They can perform overall control, but it is difficult to achieve fine-grained control over each element of the video. Additionally, training an I2V generation model with high alignment between motion information in the text and the video relies heavily on the quality of training data [4, 67, 55, 8, 50, 54], and obtaining high-quality training data is also labor-intensive.
@@
As illustrated in Figure 1(b), some controllable I2V generation models adopt well-designed modules for controlling different motion types, such as object movement [58, 69, 60, 77, 75, 39, 59, 57] or camera motion [20, 64, 65, 56]. Some methods are proposed for a single control condition, making it difficult to control multiple motion types simultaneously. Additionally, inputting these control conditions (such as point trajectories or camera extrinsics) may require expert knowledge, which creates a barrier for non-professional users. These methods are less flexible than those that directly control video motion through text input.
@@
To this end, we propose MotionAgent, which enables fine-grained control for text-guided I2V generation. The key design is the motion field agent that converts the motion information in the text prompts into explicit motion representations, providing precise motion guidance. As shown in Figure 1(c), our framework directly takes the text as input. First, the agent parses and converts the motion information in the text into object trajectories and camera extrinsics, which explicitly represent object movement and camera motion, respectively. Then, these two intermediate representations are integrated in the 3D space and project into unified optical flow maps through the proposed analytical optical flow composition module. Finally, we tune an optical flow adapter, using the unified flow as conditions, to control a base I2V diffusion model for video generation. Additionally, we design a rethinking mechanism, which forms a feedback loop enabling the agent to correct and improve the former object trajectories and camera extrinsics according to the generated video.
@@
In the experiments, we first evaluate our method on a public I2V generation benchmark [26]. Evaluation results demonstrate that our approach significantly improves the control accuracy of camera motion and achieves comparable video quality with other advanced models. Furthermore, we construct a subset from VBench for assessing the alignment of motion information in the text and the generated video. The results illustrate that our method achieves the best motion accuracy. We also conduct a user study, showing that our generated videos align better with the motion information in the text and maintain high quality. Finally, we conduct ablation studies to verify the effectiveness of the proposed components. And, we demonstrate the robustness and generalization of our method. Our contributions can be summarized as follows:
@@ 2.1 Video Generation via Agent
A series of studies propose to exploit the knowledge of agents for achieving controllable generation [32, 14, 33], zero-shot generation [25, 35, 24, 40], or long video generation [78, 53]. Anim-Director [30] builds an autonomous animation-making agent for generating contextually coherent animation videos. DreamFactory [61] leverages multi-agent collaboration principles and a key frames iteration method to ensure consistency and style across long videos. VIDEOAGENT [48] aims to improve video plan generation based on a feedback pipeline. ChatCam [34] introduces a system that navigates camera motion through conversations with users and mimics a professional cinematographer’s workflow. The most similar work is Motionzero [49], which utilizes LLM to parse the text prompt and warps the intermediate feature maps according to the extracted motion priors to achieve zero-shot video generation.
@@ 2.2 Text-guided I2V Generation
Generally, most I2V generation models [11, 73, 36, 27] support text and image input simultaneously. The input image guides the visual content of the video, while the text indicates the potential motion. VideoCrafter [7], Dynamicrafter [63], ConsistI2V [44], and I2VGen-XL [74] are all based on the U-Net [45] backbone and use a text encoder to inject text embeddings into visual features, enabling text control over video content. CogVideoX [68] and HunyuanVideo [28] based on the DIT [42] architecture, gain control capability by jointly learning text tokens and visual tokens. These methods, which directly learn the similarity between text embeddings and visual features, often highly depend on the quality of training data. However, there is a lack of large high-quality datasets with text labels that describe video motion. The text labels in most video datasets [4, 67, 55, 8, 50, 54] describe the content of a frame without much motion information and are not suitable for training I2V generation models with strong motion control capability, which leads to poor performance of these methods in achieving precise motion control through direct text input.
@@ 2.3 I2V Generation via Motion Control Module
With the development of video generation models, controllable video generation gradually attracts more attention. Existing methods [75, 64, 65] propose designing specialized modules for certain motion types, such as object movement or camera motion. DragNUWA [69] encodes sparse point trajectories into dense features as guidance information, which is then injected into the diffusion model to control object motion. DragAnything [60] uses masks to identify a central point and subsequently generates a Gaussian map to track this center, providing a guiding trajectory for object-controllable I2V generation. TrackGo [77] introduces a TrackAdapter, which encodes object trajectories into the network through a dual attention mechanism to control object movement in the generated videos. ObjCtrl-2.5D [57] extends 2D trajectory with depth and uses camera pose to represent the 2.5D object movement. Other video generation methods achieve control of camera motion. MotionCtrl [58] directly encodes camera extrinsics as features and inputs them to the model. IMAGE CONDUCTOR [31] and CameraCtrl [20] both convert sparse camera extrinsics into dense feature maps using Plücker embeddings and employ these dense embeddings to control camera motion. CPA [56] deploy the sparse motion encoding module to transform camera pose into a spatial-temporal embedding for camera control. Most of these methods can only control a single motion type at a time and struggle to achieve simultaneous control of all types.
In addition to encoding a single motion condition into deep features, some methods generate explicit intermediate representations to control multiple motion types of the generated video. Motion-I2V [46] generates optical flow maps as the intermediate representation using a flow diffusion model. MOFA-Video [39] proposes a generative motion field adapter, converting sparse point trajectories into dense optical flow through a sparse-to-dense network. AnimateAnything [29] designs a unified flow generation module, which generates optical flow maps from point trajectories and camera paths. I2VControl [13] and Motion Prompting [16] lift image to 3D and employ point trajectories as explicit intermediate representations for controlling. Diffusion as Shader [17], Perception-as-Control [10] and MotionCanvas [62] also lift input image into 3D space and utilize different intermediate motion representations for controllable video generation, such as colorize 3D points, UV textures and 3D Bounding-box. Although these models achieve control on various motion types, they often require specific knowledge, such as providing object trajectories or camera extrinsics, which creates a barrier for users.
Different from previous works, our approach proposes a motion field agent that allows direct control of the generated video through text input rather than manual motion representing input. And, we introduce an analytical optical flow composition module, which lifts the image to 3D space and enables our method to simultaneously control both object movement and camera motion.
@@ 3 Method
Figure 2: Pipeline of Motion Field Agent. Step 1: the agent first parses the input text, dividing the motion information into two parts that respectively describe object movement and camera motion. Step 2: the agent draws the object trajectories according to the text of object movement. Step 3: the agent directly generates the camera extrinsics based on the text of camera motion. Step 4 (optional): the agent rethinks and corrects the former actions according to the generated video.
Our method aims to achieve fine-grained controllable I2V generation through text input, allowing precise control of object movement and camera motion. Specifically, a motion field agent first converts the motion information in the text into object trajectories and camera extrinsics. These two explicit intermediate representations are then fed into an improved controllable I2V generation model to guide the video generation. Within the controllable I2V generation model, an analytical optical flow composition module integrates the object trajectories and camera extrinsics to compute unified optical flow maps. Subsequently, we employ Stable Video Diffusion (SVD) [5] as the base I2V diffusion model and utilize an optical flow adapter [39] as the motion control module to generate the final video.
@@ 3.1 Motion Field Agent
We generally decompose a video’s motion field into object movement within the frame and overall camera motion. Based on this decomposition, our agent performs two key tasks: object trajectory plotting and camera extrinsic generation. We also design a rethinking mechanism, which allows the agent to correct the former actions according to the generated video. The details of the designed agent can be found in Figure 2.
@@ 3.1.2 Object Trajectory Plotting
Trajectory Plotting: (4) The agent locates the object in the image based on the detection and segmentation results, thereby determining the starting point of the trajectory. (5) The agent plots the trajectory with varying lengths and curvature according to the complexity of the object movement.
To enhance the capabilities of trajectory plotting, we replace direct trajectory generation with a grid selection approach [72]. An image is divided into $N\times M$ grids, each labeled by an integer. The agent determines each point of the trajectory by selecting grid numbers, and the trajectory is formed by sequentially connecting all the points. More details of trajectory plotting can be found in the supplementary material.
@@ 3.1.3 Camera Extrinsics Generation
Agent directly generates camera extrinsics $E$ based on the text of camera motion and input image. The text description specifies the camera’s path, and the image helps the agent determine the magnitude of motion. For instance, a wide landscape scene may require intensive camera motion, while a narrow close-up shot may only need slight adjustments of camera location. We constrain the translation $T$ of the camera extrinsics to (1,1)(-1,1). In the next optical flow composition module, we rescale the translation $T$ back to a reasonable range based on the estimated depth map.
3.1.4 Rethinking
We introduce an optional rethinking mechanism for the agent to correct the former actions based on the already generated results. Specifically, the agent analyzes the generated video and reviews the actions it made at each step previously, which forms a feedback loop and enables the agent to examine and correct the former actions according to the misalignment of text prompt and generated video.
@@ 3.2.1 Analytical Optical Flow Composition
We first lift the input image to 3D space by using Metric3D [70] to estimate the depth map $D$ of the input image and unproject each pixel $I^{0}$ in the image to obtain their 3D locations $P^{0}$. CMP [71] is utilized to estimate the dense optical flow of object movement $F_{obj}$ from the generated object trajectories. Based on the estimated optical flow map $F_{obj}$ and the depth map $D$, we compute the 3D location offsets $O$ raised by object movement. Then, we add these offsets $O$ to the initial 3D locations $P^{0}$ and get the moved 3D locations ${P}^{1}$.
Subsequently, we reproject these new 3D locations $P^{1}$ into the corresponding image coordinate systems according to the generated camera extrinsics $E$ and calculate the new pixel positions $I^{1}$ in the corresponding image coordinate systems, which can be calculated as,
$I^{1}=\Pi(EP^{1}),$ (1)
where, $E$ is the generated camera extrinsics, and $\Pi$ is the projection operation. We assume that the camera coordinate of the first frame serves as the world coordinate, and we scale the translation $T$ in the generated camera extrinsics $E$ according to the maximum depth range.
Finally, we compute the offsets of the corresponding pixels in the image coordinate systems as the unified optical flow, which can be noted as,
$F=I^{1}-I^{0},$ (2)
where $F$ are named unified optical flow maps. These flow maps contain motion information of both object movement and camera motion.
@@ 3.2.2 Optical Flow Adapter Tuning
Since the unified optical flow maps contain geometrically unrealistic regions compared to the true optical flow maps, directly using these flow maps as the input of the optical flow adapter, which is trained on real optical flow maps, may lead to degraded results in the generated videos. So, we propose fine-tuning the optical flow adapter on the unified optical flow maps.
Specifically, we first use an optical flow model, Unimatch [66], to estimate the real optical flow. Then we employ an SLAM method, DROID-SLAM [52], to compute the camera extrinsics for each frame. Next, we eliminate the optical flow caused by camera motion based on the estimated camera extrinsics, and approximately obtain the optical flow only caused by object movement. We perform sparse sampling [71] on these optical flow maps to obtain sparse object trajectories. More details of data preparation can be found in the supplementary material. Subsequently, we reuse the proposed analytical composition method to calculate the unified optical flow maps utilized as input to fine-tune the optical flow adapter. This step eliminates the domain gap between real and unified flow fields and facilitates high-quality video results.
@@ 4.1 Implementation Details
Metrics. In the comparison experiments of the general I2V generation task, we adopt the same evaluation metrics introduced by VBench [26]. In the comparison of controllable I2V generation, we report Object Movement Q&A, Complex Camera Motion, and Overall Score. All evaluation metrics are higher-the-better. The details of evaluation metrics can be found in the supplementary materials.